OpenStudy (anonymous):

Will give medal and fan Use the equation and type the ordered-pairs. y = 2^x {(-1,__ a0), (0,__ a1), (1,__ a2), (2,__ a3), (3,__ a4), (4,__ a5)}

OpenStudy (johnweldon1993):

Alright...so this is saying to make the ordered pairs using \[\large y = 2^x\] The first one (a0 give (-1 , ?) so we take that -1 and plug it in for 'x' in this equation...and see what the ? would be \[\large y = 2^{-1} = \frac{1}{2} = .5\] So the first ordered pair a0 will be (-1 , .5)

OpenStudy (johnweldon1993):

and so on and so forth... so a1 would then be what would the equation be when x = 0 \[\large y = 2^0 = 1\] so that ordered pair a1 would be (0 , 1)

OpenStudy (anonymous):

ohhhhh alright yeah i see now thank you!

OpenStudy (johnweldon1993):

Anytime :)